# Denoising Diffusion Variational Inference: Diffusion Models as Expressive Variational Posteriors

### v4

@@ Appendix C Comparision of Methods
We provide a comprehensive comparison of different methods in Table 4. Vahdat, Kreis, and Kautz (2021); Wehenkel and Louppe (2021); Rombach et al. (2022) perform diffusion in the latent space of a VAE to improve the efficiency of image generation. Their goal is high sample quality, and they introduce into $p$ hierarchical latents with simple Gaussian priors. Our goal is different: we seek a method to fit a $p$ with structured latents (e.g., in probabilistic programming or in science applications, users introduce prior knowledge via hand-crafted $p$), and we improve variational inference in this structured model by introducing auxiliary latents into $q$.
Recent work (Preechakul et al. 2022; Zhang, Zhao, and Lin 2022; Wang et al. 2023) has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$. Cohen et al. (2022) crafts a diffusion bridge linking a continuous coded vector to a non-informative prior distribution.
@@ Appendix E Connections to Diffusion Samplers
Diffusion sampling (Berner, Richter, and Ullrich 2022; Zhang and Chen 2021; Vargas, Grathwohl, and Doucet 2023; Zhang et al. 2023; Richter, Berner, and Liu 2023; Sendera et al. 2024; Akhound-Sadegh et al. 2024) mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function. These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift), leading to several approaches, e.g., path integral sampler (PIS) (Zhang and Chen 2021), denoising diffusion sampler (DDS) (Vargas, Grathwohl, and Doucet 2023), and time-reversed diffusion sampler (DIS) (Berner, Richter, and Ullrich 2022), which have been unified by Richter, Berner, and Liu (2023). Some other works (Zhang et al. 2023; Akhound-Sadegh et al. 2024) use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration. Sendera et al. (2024) benchmarked these previous diffusion-structured amortized inference methods and studied how to improve credit assignment in diffusion samplers, which refers to the propagation of learning signals from the target density to the parameters of earlier sampling steps. Overall, there are indeed some strong connections between these works and ours:
@@
  • To tackle the challenge of credit assignment – propagating weak learning signals through the sampling trajectory, the techniques proposed in diffusion-structured samplers are mostly based on partial trajectory information, which has higher training costs over on-policy (Zhang and Chen 2021) or off-policy (Malkin et al. 2022) trajectory-based optimization. Instead, we introduce a wake-sleep optimization algorithm and its simplified version to alleviate the weak learning signal issue and optimize the evidence lower bound in a better way.
@@
(Berner, Richter, and Ullrich 2022; Zhang and Chen 2021; Vargas, Grathwohl, and Doucet 2023; Akhound-Sadegh et al. 2024) (Zhang et al. 2023; Richter, Berner, and Liu 2023; Sendera et al. 2024; Malkin et al. 2022)
@@ Appendix F Priors
Pinwheel. This distribution was used in (Johnson et al. 2016). We define the number of clusters to be 10. For semi-supervised learning experiments, this prior is partitioned into 10 partitions, each partition being a cluster.
Swiss Roll. This distribution was used in (Marsland 2014). For semi-supervised learning experiments, this prior is partitioned into 10 partitions. The samples from the prior can actually be characterized by a single scalar representing how far you are long the swiss roll from the center. The paritioning is done by creating 10 equal-length intervals in this 1D space.
@@ Appendix I Genotype Analysis Experiments Details
AEVB $-12.96\pm 1.81$ $3.26\pm 0.60$ $-12.87\pm 4.55$ $6.25\pm 1.58$ $-7.91\pm 0.11$ $2.91\pm 0.17$
AEVB-IAF $-3.24\pm 0.16$ $1.71\pm 0.84$ $-4.03\pm 0.73$ $5.51\pm 0.51$ $\mathbf{-2.10\pm 0.31}$ $1.71\pm 0.77$
@@
H-AEVB $-4.42\pm 0.46$ $\mathbf{1.69\pm 0.17}$ $-5.36\pm 0.77$ $5.74\pm 0.55$ $-2.86\pm 0.11$ $1.64\pm 0.09$
PIS ${-2.92\pm 1.23}$ ${3.61\pm 0.62}$ $-4.14\pm 0.49$ $7.14\pm 0.14$ $-4.85\pm 0.06$ $3.91\pm 0.06$
DDVI $\mathbf{-1.38\pm 0.44}$ ${1.75\pm 0.53}$ $\mathbf{-3.05\pm 0.65}$ ${5.66\pm 2.63}$ $-2.47\pm 0.30$ $\mathbf{1.58\pm 0.09}$
@@
Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $-17.14\pm 1.46$ $0.30\pm 0.05$ $2.32\pm 0.27$ $-17.89\pm 5.21$ $0.20\pm 0.07$ $6.56\pm 2.25$ $-13.30\pm 1.50$ $0.30\pm 0.05$ $1.95\pm 0.28$ AEVB-IAF $-5.70\pm 0.07$ $0.47\pm 0.01$ $1.62\pm 0.05$ $-5.53\pm 2.82$ $0.28\pm 0.08$ $6.82\pm 1.90$ $-4.41\pm 0.53$ $0.36\pm 0.01$ $1.58\pm 0.15$ AAEB $N/A$ $0.25\pm 0.01$ $1.77\pm 0.14$ $N/A$ $0.23\pm 0.01$ $3.38\pm 0.30$ $N/A$ $0.23\pm 0.04$ $1.74\pm 0.15$ DDVI $\mathbf{-1.60\pm 0.29}$ $\mathbf{0.49\pm 0.01}$ $\mathbf{1.09\pm 0.05}$ $\mathbf{-4.13\pm 1.51}$ $\mathbf{0.47\pm 0.09}$ $\mathbf{2.29\pm 0.08}$ $\mathbf{-1.73\pm 0.64}$ $\mathbf{0.49\pm 0.01}$ $\mathbf{1.48\pm 0.02}$
@@
Method Latent NLL - Pinwheel Latent NLL - Swiss Roll Latent NLL - Square AEVB $1.68\pm 0.31$ $5.65\pm 1.58$ $2.78\pm 0.61$ AEVB-IAF $1.64\pm 0.73$ $4.43\pm 1.09$ $1.68\pm 0.27$ AAEB -- -- -- H-AEVB $2.25\pm 3.02$ $4.04\pm 4.62$ $2.22\pm 2.03$ DDVI $\mathbf{1.27\pm 0.21}$ $\mathbf{3.86\pm 1.17}$ $\mathbf{1.56\pm 0.09}$ DDVI (w/o sleep term) $2.12$ $5.25$ $2.97$
@@
[…]
@@
This form mirrors the ELBO derivation for diffusion models (Sohl-Dickstein et al. 2015), where each step in the Markov chain contributes a KL divergence term, and the reconstruction term arises from the connection between the noisy latent variables and the original data.

### v5

@@ Appendix C Comparision of Methods
We provide a comprehensive comparison of different methods in Table 4. vahdat2021score; wehenkel2021diffusion; rombach2022high perform diffusion in the latent space of a VAE to improve the efficiency of image generation. Their goal is high sample quality, and they introduce into $p$ hierarchical latents with simple Gaussian priors. Our goal is different: we seek a method to fit a $p$ with structured latents (e.g., in probabilistic programming or in science applications, users introduce prior knowledge via hand-crafted $p$), and we improve variational inference in this structured model by introducing auxiliary latents into $q$.
Recent work (preechakul2022diffusion; zhang2022unsupervised; wang2023infodiffusion) has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$. cohen2022diffusion crafts a diffusion bridge linking a continuous coded vector to a non-informative prior distribution.
@@ Appendix E Connections to Diffusion Samplers
Diffusion sampling (berner2022optimal; zhang2021path; vargas2023denoising; zhang2023diffusion; richter2023improved; sendera2024diffusion; akhound2024iterated) mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function. These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift), leading to several approaches, e.g., path integral sampler (PIS) (zhang2021path), denoising diffusion sampler (DDS) (vargas2023denoising), and time-reversed diffusion sampler (DIS) (berner2022optimal), which have been unified by richter2023improved. Some other works (zhang2023diffusion; akhound2024iterated) use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration. sendera2024diffusion benchmarked these previous diffusion-structured amortized inference methods and studied how to improve credit assignment in diffusion samplers, which refers to the propagation of learning signals from the target density to the parameters of earlier sampling steps. Overall, there are indeed some strong connections between these works and ours:
@@
  • To tackle the challenge of credit assignment – propagating weak learning signals through the sampling trajectory, the techniques proposed in diffusion-structured samplers are mostly based on partial trajectory information, which has higher training costs over on-policy (zhang2021path) or off-policy (malkin2022trajectory) trajectory-based optimization. Instead, we introduce a wake-sleep optimization algorithm and its simplified version to alleviate the weak learning signal issue and optimize the evidence lower bound in a better way.
@@
(berner2022optimal; zhang2021path; vargas2023denoising; akhound2024iterated) (zhang2023diffusion; richter2023improved; sendera2024diffusion; malkin2022trajectory)
@@ Appendix F Priors
Pinwheel. This distribution was used in (johnson2016svae). We define the number of clusters to be 10. For semi-supervised learning experiments, this prior is partitioned into 10 partitions, each partition being a cluster.
Swiss Roll. This distribution was used in (marsland2014swissroll). For semi-supervised learning experiments, this prior is partitioned into 10 partitions. The samples from the prior can actually be characterized by a single scalar representing how far you are long the swiss roll from the center. The paritioning is done by creating 10 equal-length intervals in this 1D space.
@@ Appendix I Genotype Analysis Experiments Details
AEVB $12.96\pm 1.81$ $3.26\pm 0.60$ $12.87\pm 4.55$ $6.25\pm 1.58$ $7.91\pm 0.11$ $2.91\pm 0.17$
AEVB-IAF $3.24\pm 0.16$ $1.71\pm 0.84$ $4.03\pm 0.73$ $5.51\pm 0.51$ $\mathbf{-2.10\pm 0.31}$ $1.71\pm 0.77$
@@
H-AEVB $4.42\pm 0.46$ $\mathbf{1.69\pm 0.17}$ $5.36\pm 0.77$ $5.74\pm 0.55$ $2.86\pm 0.11$ $1.64\pm 0.09$
PIS $2.92\pm 1.23$} $3.61\pm 0.62$} $4.14\pm 0.49$ $7.14\pm 0.14$ $4.85\pm 0.06$ $3.91\pm 0.06$
DDVI $\mathbf{-1.38\pm 0.44}$ $1.75\pm 0.53$} $\mathbf{-3.05\pm 0.65}$ $5.66\pm 2.63$} $2.47\pm 0.30$ $\mathbf{1.58\pm 0.09}$
@@
Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $17.14\pm 1.46$ $0.30\pm 0.05$ $2.32\pm 0.27$ $17.89\pm 5.21$ $0.20\pm 0.07$ $6.56\pm 2.25$ $13.30\pm 1.50$ $0.30\pm 0.05$ $1.95\pm 0.28$ AEVB-IAF $5.70\pm 0.07$ $0.47\pm 0.01$ $1.62\pm 0.05$ $5.53\pm 2.82$ $0.28\pm 0.08$ $6.82\pm 1.90$ $4.41\pm 0.53$ $0.36\pm 0.01$ $1.58\pm 0.15$ AAEB $N/A$ $0.25\pm 0.01$ $1.77\pm 0.14$ $N/A$ $0.23\pm 0.01$ $3.38\pm 0.30$ $N/A$ $0.23\pm 0.04$ $1.74\pm 0.15$ DDVI $\mathbf{-1.60\pm 0.29}$ $\mathbf{0.49\pm 0.01}$ $\mathbf{1.09\pm 0.05}$ $\mathbf{-4.13\pm 1.51}$ $\mathbf{0.47\pm 0.09}$ $\mathbf{2.29\pm 0.08}$ $\mathbf{-1.73\pm 0.64}$ $\mathbf{0.49\pm 0.01}$ $\mathbf{1.48\pm 0.02}$
@@
Method Latent NLL - Pinwheel Latent NLL - Swiss Roll Latent NLL - Square AEVB $1.68\pm 0.31$ $5.65\pm 1.58$ $2.78\pm 0.61$ AEVB-IAF $1.64\pm 0.73$ $4.43\pm 1.09$ $1.68\pm 0.27$ AAEB - - - H-AEVB $2.25\pm 3.02$ $4.04\pm 4.62$ $2.22\pm 2.03$ DDVI $\mathbf{1.27\pm 0.21}$ $\mathbf{3.86\pm 1.17}$ $\mathbf{1.56\pm 0.09}$ DDVI (w/o sleep term) $2.12$ $5.25$ $2.97$
@@
$\log p_{\boldsymbol{\theta}}(\mathbf{x})$ $=\log\int p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})\,d\mathbf{z}$ (15)
$\geq\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log\frac{p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})}{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\right]\quad\text{(by Jensen's inequality)}$ (16)
@@
This form mirrors the ELBO derivation for diffusion models (sohl2015deep), where each step in the Markov chain contributes a KL divergence term, and the reconstruction term arises from the connection between the noisy latent variables and the original data.